Cloud At Cost
   

Client portal

  • Network Status Update (Resolved)

    May 1, 2015 by Jason Barnes
  • ### 5/1/2015 3:30PM EST

    The Rogers Tech team has advised that the circuit is back up, stable and repair work is complete.  You should no longer see any flaps. They are still waiting on technician notes on the work completed and more updates will follow on the full cause and resolution of this incident. 


    ### 5/1/2015 11:30AM EST

    CloudatCost is working at moving all Dev/BigDog Servers to Redundant links used by CAC Enterprise and CloudPro Services.


    ### 5/1/2015 11AM EST

    Rogers Update

    We are pending the arrival of our fibre crews to the cut location. In parallel, we are investigating if traffic can be rerouted around the damage.



    ### 5/1/2015 10AM EST


    Here is the latest update from Rogers Tech.


    "The latest update is that our head end crews have identified another fibre break, separate from the one yesterday morning. We have a fibre crew currently on-route, but cannot provide an ETR"

    We will continue to update this thread.


    To Unsubscribe : If you don't want to receive these messages please change your email preferences here https://members.cloudatcost.com on your profile page.


Powered by